OpenStudy (abhisar):

Given that magnitude of A= magnitude of B. What is the angle between (A+B) and (A-B) ?

OpenStudy (abhisar):

@hartnn

OpenStudy (abhisar):

Sorry m big noob in mathematics

OpenStudy (abhisar):

Ok i know geometrically, it will be 90°. Can u help me finding in a numerical way ?

hartnn (hartnn):

yes know the formula for dot product ?

OpenStudy (abhisar):

yep

OpenStudy (abhisar):

A.B= ABCos\(\theta\)

hartnn (hartnn):

so, cos theta = x.y /(|x||y|) right ?? i took x and y because A,B would be confusing x= A+B, y =A-B

OpenStudy (abhisar):

ok..

hartnn (hartnn):

just try to find the numerator of cos theta (A+B). (A-B) = .... ?

OpenStudy (abhisar):

A\(^2\) + B\(^2\) + 2ABCos\(\theta\) ..... like this ?

hartnn (hartnn):

no... foil it (A+B).(A-B) = ... ? (like z(x+y) = xz +yz)

hartnn (hartnn):

first term will be \(A.A = A (dot)A = |A|^2 \)

hartnn (hartnn):

don't forget the dots...its a dot product operation A.B -B.A = A.B -A.B = 0 (because dot product is commutitive)

OpenStudy (abhisar):

oh..yes..i got it !

hartnn (hartnn):

so, numerator is |A|^2 -|B|^2 and it is given that |A| = |B| when is cos theta = 0 ?

OpenStudy (abhisar):

90

OpenStudy (abhisar):

great !

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (abhisar):

Thnx once again !! \(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)